OpenStudy (mb194227):

Please help me find all the real zeroes of the following function: g(x) = 2x^3 + x^2 - 6x - 3.

satellite73 (satellite73):

factor out the common factor of \(x\) first, then you will have a quadratic to solve

satellite73 (satellite73):

ok skip that, i didn't see the 3 at the end it is \[ g(x) = 2x^3 + x^2 - 6x - 3\] right?

OpenStudy (mb194227):

Correct.

satellite73 (satellite73):

try some easy numbers that divide 3 evenly first 1 doesn't work how about \(-1\)?

satellite73 (satellite73):

nope that doesn't work either dang

satellite73 (satellite73):

neither does 3 or -3, on to the fractions

satellite73 (satellite73):

try \(x=-\frac{1}{2}\) i think that one works

OpenStudy (mb194227):

Well, I know the steps, but I need help executing them: 1. Find all possible rational roots; the factor of the constant / the factor of the leading coefficient (p/q). 2. Test all possible roots of p/q via synthetic division. 3. Once you've found a quadratic equation, solve it for the other roots. However, I do not know how to use p/q. May you teach me how?

OpenStudy (mb194227):

Yes, x = -1/2 works.

satellite73 (satellite73):

the constant is -3 the leading coefficient is 2 the \(\frac{p}{q}\) business is fraction (or integers) where \(p\) divides the constant, and \{q\) divides the leading coefficient

satellite73 (satellite73):

number that go in to 3 evenly are \[\pm1,\pm3\] that is why i suggested trying those first, because they are easiest

satellite73 (satellite73):

but none of those worked so now the other possible rational zeros are \[\pm\frac{1}{2}, \pm\frac{3}{2}\]

satellite73 (satellite73):

those are the numbers where the numerator divides 3 and the denominator divides 2

OpenStudy (mb194227):

Okay, thank you! I finally have a better understanding of solving polynomials.

satellite73 (satellite73):

yw